# Evolution and compression in LLMs: On the emergence of human-aligned categorization

### v3

@@ Appendix B English color naming procedure
The color naming task was conducted by prompting a model to label a color chip, for all $330$ color chips from the WCS stimulus array, one stimulus at a time, in a randomized order. Unlike the prompts in our IICLL study, we do not provide the model with access to previous interactions. Each prompt was of the form: “What color is this [ <color coordinates> ]? You may only use one of the following allowed labels: [‘Red’, ‘Blue’, $\dots$] Please provide only a single label from the list just provided. Do not give any explanation.” When prompting models with text only, the color coordinates were sRGB triples. The allowed set of labels were fourteen basic English color terms corresponding to the modal terms from Lindsey2014Color.
@@ Appendix C English color naming with CIELAB inputs
Figure 6 illustrates mode maps of the resulting systems for models when prompted with CIELAB-encoded colors. We found that performance generally degraded across all models when providing features as CIELAB triples: they struggled significantly to align with established English color boundaries, and the resulting category systems were considerably noisier compared to the sRGB-based outputs. This outcome is consistent with previous research by Marjieh2024Large which noted that other frontier models failed to produce coherent English color category systems when prompted with CIELAB triples. Based on this observation, we restricted our IICLL study to the sRGB feature space to assess the models’ inductive bias under ideal conditions.
@@ Appendix E Naming systems for all models
Figure 7: IB complexity-accuracy tradeoffs achieved by text-based LLMs. Same as Figure 2a but including base (not instruction-tuned) models.
Figure 8: IB complexity-accuracy tradeoffs achieved by instruction-tuned multimodal LLMs. Points without color filling indicate systems resulting from providing colors as images, rather than via sRGB coordinates in text. Each minimal pair is connected by a line. Interestingly, there appears to be a cutoff around 3 bits of complexity for which models that already perform well with text-based prompting do not improve alignment to English when prompted with an image instead.
@@
As can be seen in Figure 9, many of the models performed poorly on the English naming task, with many base (not instruction-tuned) models and some smaller instruction-tuned models failing to produce naming systems with any coherent category structure. Figure 7 shows that these systems are distributed widely across the information plane, with only larger, instruction-tuned, very recent models achieving similar IB-efficiency tradeoffs to English speakers. Interestingly, some models show similar category structure not to English, but to languages in the WCS. For example, Olmo-2-32B-Instruct, and Qwen-2.5-VL-Instruct (when presented with images) bear resemblance to Mayoruna (see the final row of Figure 13), as well as some of the systems to which Gemini-2.0 converges over generations of IICLL. This suggests that although some larger, instruction tuned models and multimodal instruction-tuned models fail to recover the English naming system, they may still possess some general bias towards color category structure that is aligned with humans.
@@ Appendix F Olmo training trajectory
As model checkpoints are available for the base Olmo 2 32B (OLMoTeam20252) model throughout its training, we repeated our English color naming study throughout different stages of its model training. The training procedure involved two main phases: pre-training on up to 6 trillion tokens, and a second training stage using a curated dataset of 843 billion tokens, including high-quality, academic, and instruction-tuning data. This also involved training on 100 billion and 300 billion token samples and then averaging the final checkpoints in a process called model souping.
Furthermore, we found that hinting to the model that the coordinates were in fact sRGB values could improve model performance. Specifically, when the prompt was changed to “What color is this rgb=[ <color coordinates> ]?”. To aid in our analysis of the base pretrained Olmo models, we used this form of prompting in order to see what structure is achievable in principle.
@@ Appendix G Iterated in-context language learning procedure
We aimed to replicate the procedure of Xu2013Cultural. The initial training data for the first generation of IICLL consisted of random pairings of colors and nonsense terms, with a sample size six times the allowed vocabulary size. For vocabulary sizes ranging from 2 to 6 terms, the initial training data was sourced directly from Xu2013Cultural to facilitate direct comparison. A 14-term condition was also included for comparison with English (Lindsey2014Color), with its initial mappings randomly generated as it was not part of the original study.
@@
Once labels were obtained for all stimuli in the production phase, this full set of stimulus-label pairs constituted the current generation’s complete color category system. This marked the completion of one generation. To initiate the next generation, a new training set was created by randomly sampling stimulus-word pairs from the just-completed generation’s system, with the sample size being six times the number of allowed words in the vocabulary. This entire process was repeated for up to 13 generations, (which was the length of chains in Xu2013Cultural) or until two successive generations yielded degenerate systems (where all stimuli are mapped to a single term).
@@ Appendix H Rotation analysis
Figure 11: Rotation analysis for final systems of all models’ IICLL chains and human ILL chains from Xu2013Cultural, WCS languages (Kay2009World). The $x$-axis denotes rotation along columns (hue dimension) of the WCS grid, while the $y$-axis is the difference $(\Delta)$ in mean (a) efficiency loss, (b) alignment to IB optima, or (c) alignment to languages in the WCS. Each colored curve is the average across initializations and conditions, and the colored region corresponds to the 95% confidence interval around the mean (assuming a normal distribution). In (a), $\Delta>0$ means that the actual systems are more efficient that their rotated counterparts, whereas in (b) and (c) $\Delta<0$ means that the actual systems are more aligned.
Figure 12: Rotation examples. Hypothetical variants for a final generation system of IICLL with Gemini 2.0 in the $k=14$ condition, and the English system from Lindsey2014Color. Variants are obtained by rotating the color naming system in the hue dimension across the columns of the WCS stimulus palette. $r=0$ corresponds to the actual system, $r=5$ corresponds to a shift of five columns to the right, and r=5r=-5 corresponds to a shift of five columns to the left.
@@ Appendix I Example chains compared to World Color Survey languages
During the production phase of our IICLL experiments, each response from the model was saved and added to a sliding window of N previous user-model interactions, excluding the initial training data. After preliminary explorations with window sizes of 0, 10, 20, and 50, we determined that a window size of 0 led to degenerate category systems more frequently, while the results from a window size of 20 and 50 did not significantly differ from those obtained with a window size of 10. Consequently, we set the window size to 10 for all IICLL experiments. This window was included in the prompt for subsequent stimuli, presented after the initial training data and before each new stimulus, aiming to promote coherence in the model’s responses and to crudely mimic the influence of short-term memory that human participants would possess in Xu2013Cultural’s experiments.
@@
We restricted our main Iterated In-Context Language Learning (IICLL) analysis to models that performed well on the initial English naming task to focus specifically on the evolution towards non-trivial IB-efficiency. This prioritization was based on our suspicion that models lacking sufficient inference power, coherence in color naming, or in-context learning capacity would be unable to sustain non-trivial category structure under IL pressure.
In order to confirm this empirically for models with weak initial performance, we ran supplementary IICLL simulations using four models that performed modestly at the English naming task: Gemma 3 4b, Llama 3.2 3B, Qwen 2.5 3B, and OLMo 2 7B (all instruction-tuned). These simulations covered the five word-number conditions ($k=2$ through $6$) used in the experiments of Xu2013Cultural.

### v4

@@ Appendix B English color naming procedure
The color naming task was conducted by prompting a model to label a color chip, for all $330$ color chips from the WCS stimulus array, one stimulus at a time, in a randomized order. Unlike the prompts in our IICLL study, we do not provide the model with access to previous interactions. Each prompt was of the form: “What color is this [ <color coordinates> ]? You may only use one of the following allowed labels: [‘Red’, ‘Blue’, $\dots$] Please provide only a single label from the list just provided. Do not give any explanation.” When prompting models with text only, the color coordinates were sRGB triples. The allowed set of labels were fourteen basic English color terms corresponding to the modal terms from Lindsey and Brown (2014).
@@ Appendix C English color naming with CIELAB inputs
Figure 6 illustrates mode maps of the resulting systems for models when prompted with CIELAB-encoded colors. We found that performance generally degraded across all models when providing features as CIELAB triples: they struggled significantly to align with established English color boundaries, and the resulting category systems were considerably noisier compared to the sRGB-based outputs. This outcome is consistent with previous research by Marjieh et al. (2024) which noted that other frontier models failed to produce coherent English color category systems when prompted with CIELAB triples. Based on this observation, we restricted our IICLL study to the sRGB feature space to assess the models’ inductive bias under ideal conditions.
@@ Appendix E Naming systems for all models
Figure 7: IB complexity-accuracy tradeoffs achieved by text-based LLMs. Same as Figure 2a but including base (not instruction-tuned) models. Points with a black edge indicate models that are instruction-tuned.
Figure 8: IB complexity-accuracy tradeoffs achieved by instruction-tuned multimodal LLMs. Points without color filling indicate systems resulting from providing colors as images, rather than via sRGB coordinates in text. Each minimal pair is connected by a line. Interestingly, there appears to be a cutoff at around 3 bits of complexity for which models that already perform well with text-based prompting do not improve alignment to English when prompted with an image instead.
@@
As can be seen in Figure 9, many of the models performed poorly on the English naming task, with many base (not instruction-tuned) models and some smaller instruction-tuned models failing to produce naming systems with any coherent category structure. Figure 7 shows that these systems are distributed widely across the information plane, with only larger, instruction-tuned, very recent models achieving similar IB-efficiency tradeoffs to English speakers. Interestingly, some models show similar category structure not to English, but to languages in the WCS. For example, Olmo-2-32B-Instruct, and Qwen-2.5-VL-7B-Instruct (when presented with images) bear resemblance to Mayoruna (see the final row of Figure 13), as well as some of the systems to which Gemini-2.0 converges over generations of IICLL. This suggests that although some larger, instruction tuned models and multimodal instruction-tuned models fail to recover the English naming system, they may still possess some general bias towards color category structure that is aligned with humans.
@@ Appendix F Olmo training trajectory
As model checkpoints are available for the base Olmo 2 32B (OLMo-Team et al., 2025) model throughout its training, we repeated our English color naming study throughout different stages of its model training. The training procedure involved two main phases: pre-training on up to 6 trillion tokens, and a second training stage using a curated dataset of 843 billion tokens, including high-quality, academic, and instruction-tuning data. This also involved training on 100 billion and 300 billion token samples and then averaging the final checkpoints in a process called model souping.
Furthermore, we found that hinting to the model that the coordinates were in fact sRGB values could improve model performance—specifically, when the prompt was changed to “What color is this rgb=[ <color coordinates> ]?”. To aid in our analysis of the base pretrained Olmo models, we used this form of prompting in order to see what structure is achievable in principle.
@@ Appendix G Iterated in-context language learning procedure
We aimed to replicate the procedure of Xu et al. (2013). The initial training data for the first generation of IICLL consisted of random pairings of colors and nonsense terms, with a sample size six times the allowed vocabulary size. For vocabulary sizes ranging from 2 to 6 terms, the initial training data was sourced directly from Xu et al. (2013) to facilitate direct comparison. A 14-term condition was also included for comparison with English (Lindsey and Brown, 2014), with its initial mappings randomly generated as it was not part of the original study.
@@
Once labels were obtained for all stimuli in the production phase, this full set of stimulus-label pairs constituted the current generation’s complete color category system. This marked the completion of one generation. To initiate the next generation, a new training set was created by randomly sampling stimulus-word pairs from the just-completed generation’s system, with the sample size being six times the number of allowed words in the vocabulary. This entire process was repeated for up to 13 generations, (which was the length of chains in Xu et al. (2013)) or until two successive generations yielded degenerate systems (where all stimuli are mapped to a single term).
@@ Appendix H Rotation analysis
Figure 11: Rotation analysis for final systems of all models’ IICLL chains and human ILL chains from Xu et al. (2013). The $x$-axis denotes rotation along columns (hue dimension) of the WCS grid, while the $y$-axis is the difference $(\Delta)$ in mean (a) efficiency loss, (b) alignment to IB optima, or (c) alignment to languages in the WCS. Each colored curve is the average across initializations and conditions, and the colored region corresponds to the 95% confidence interval around the mean (assuming a normal distribution). In (a), $\Delta>0$ means that the actual systems are more efficient that their rotated counterparts, whereas in (b) and (c) $\Delta<0$ means that the actual systems are more aligned.
Figure 12: Rotation examples. Hypothetical variants for a final generation system of IICLL with Gemini 2.0 in the $k=14$ condition, and the English system from Lindsey and Brown (2014). Variants are obtained by rotating the color naming system in the hue dimension across the columns of the WCS stimulus palette. $r=0$ corresponds to the actual system, $r=5$ corresponds to a shift of five columns to the right, and r=5r=-5 corresponds to a shift of five columns to the left.
@@ Appendix I Example chains compared to World Color Survey languages
During the production phase of our IICLL experiments, each response from the model was saved and added to a sliding window of N previous user-model interactions, excluding the initial training data. After preliminary explorations with window sizes of 0, 10, 20, and 50, we determined that a window size of 0 led to degenerate category systems more frequently (see Figures 14 and 15), while the results from a window size of 20 and 50 did not significantly differ from those obtained with a window size of 10. Consequently, we set the window size to 10 for all IICLL experiments. This window was included in the prompt for subsequent stimuli, presented after the initial training data and before each new stimulus, aiming to promote coherence in the model’s responses and to crudely mimic the influence of short-term memory that human participants would possess in Xu et al. (2013)’s experiments.
@@
We restricted our main Iterated In-Context Language Learning (IICLL) analysis to models that performed well on the initial English naming task to focus specifically on the evolution towards non-trivial IB-efficiency. This prioritization was based on our expectation that models lacking sufficient inference power, coherence in color naming, or in-context learning capacity would be unable to sustain non-trivial category structure under IL pressure.
In order to confirm this empirically for models with weak initial performance, we ran supplementary IICLL simulations using four models that performed modestly at the English naming task: Gemma 3 4b, Llama 3.2 3B, Qwen 2.5 3B, and OLMo 2 7B (all instruction-tuned). These simulations covered the five vocabulary-size conditions ($k=2$ through $6$) used in the experiments of Xu et al. (2013).